Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  RH-ISAC, Page 4 of 140  Not logged in ELOG logo
ID Date Author Category Type Module Target/Number Subject Text Attachments
  2742   Tuesday, January 07, 2025, 12:05 David WangITEStandard OperationTM3UCx #47ITE primary and secondary vacuum vessel leak check for 2025 shut downITE primary and secondary vacuum vessel
leak check for 2025 shut down have been done.
No leak.  The TM3 waterlines helium
  
  2741   Monday, January 06, 2025, 08:38 Frank SongITWStandard OperationTM4UCx#46module disconnection TM4 WITH UCx#46 was disconnected
in 2025-01-02 in ITW. Forms attached.
 TM4_with_UCx#46_disconnection_in_ITW_2025-01-02.pdf 
  2740   Friday, January 03, 2025, 12:34 Frank Song5 Shielding BlocksMaintenance UCx#47shielding blocks moving Five shielding blocks are moved from
ITE to storage area ( one of them is sitting
on B2 as limit of room) for target stations
  
  2739   Friday, January 03, 2025, 10:41 David WangITWStandard OperationTM4UCX46ITW primary and secondary vacuum vessel leak check for 2025 shut downITW primary and secondary vacuum vessel
leak check for  2025 shut down has been
done. Also,TM4 internal water line helium
 ITW_primary_and_secondary_vacuum_leak_check_2025.pdf 
  2738   Friday, January 03, 2025, 09:59 Adam NewsomeFacilitiesStandard Operation  Safety Walkaround complete - ISAC Hot Cells, Target HallA safety walkaround was completed for the
ISAC Hot Cells and Target Hall Areas.
The resulting spreadsheet can be
  
  2737   Wednesday, December 18, 2024, 14:17 Adam NewsomePail Handling ToolMaintenance  Pail Handling Tool - InspectionAn informal inspection of the main Pail
Handling Tool was performed today as the
final one for the year. Note that the main
 ISAC_Pail_Handling_Tool_Inspection_Record_-_2024.pdf 
  2736   Wednesday, December 18, 2024, 12:18 Adam NewsomeSpent Target VaultStandard Operation  Five Pails Transferred from Main Vault to Mini VaultFive pails were transferred from main vault
to mini vault. Note that there was already
one pail in the mini vault, so at present
 MiniVault_2024-12-18.pdf 
  2735   Monday, December 09, 2024, 13:10 Frank Song5 Shielding BlocksStandard OperationTM3UCx#47shielding blocks transfer 5 Shielding Blocks have been transferred
from ITW to ITE. MAA is secured after moving.
  
  2734   Thursday, December 05, 2024, 12:00 Adam NewsomeSpent Target VaultStandard Operation Ta#68Spent target move - Ta#68 in Pail 300Ta#68 spent target was moved in Pail 300
to storage vault location 5A. The dose rate
measured at the South Hot Cell was 429 mSv/h
 Storage_Vault_Contents_2024-12-05.pdf 
  2733   Thursday, December 05, 2024, 11:43 Aaron TamSouth Hot-CellStandard Operation Ta#68 Pail loading - Ta#68 & Others into Pail#300December 5, 2024 - Aaron Tam w/ help from
Chad, Peter and Adam at the Crane  
Hot Cell Prep:
 Pail_Landing.jpgLoaded_Pail.jpgParasitic_Lid_that_was_missed.jpgLocation_of_all_parasitic_type_targets_and_fasteners.jpgParastics_and_Target_stuff_that_went_inside_this_pail.jpg 
  2732   Wednesday, December 04, 2024, 14:30 Aaron TamSouth Hot-CellStandard Operation Ta#68Ta#68 PIE Abstract:
This target has seen some irregular
beam patterns 
 32x 
  2731   Wednesday, December 04, 2024, 10:40 Adam NewsomeSouth Hot-CellStandard OperationTM2 TM2 move - SHC to siloTM2 was remotely moved from the South Hot
Cell to SW Silo. Reading at SHC = 89.8 mSv/h.
The move was smooth. 
  
  2730   Wednesday, December 04, 2024, 09:00 David WangFacilitiesStandard OperationTM2 all water lines on TM2 have been purged.all water lines on TM2 have been purged.
Refilled them(except HS ) with 60psi nitrogen
for module silo storage before next year
  
  2729   Wednesday, December 04, 2024, 08:05 Frank SongITEStandard OperationTM3UCx#47Module ConnectionTM3 with new target UCx#47 was connected
in ITE right after be moved from TCS. 
 TM3_with_new_target_UCx#47_connection_in_ITE_2024-12-03.pdf 
  2728   Tuesday, December 03, 2024, 15:53 Aaron TamSouth Hot-CellStandard OperationTM2Ta#68Ta#68 Removal  Note: Ta#68 had a difficult
install. See E-log ID# 2698 for more info 
First look:
 22x 
  2727   Tuesday, December 03, 2024, 12:28 Adam NewsomeSouth Hot-CellStandard OperationTM2 TM3 move - TCS to ITETM3 was moved with UCx#47 from TCS to ITE.
The move was smooth. 
  
  2726   Tuesday, December 03, 2024, 10:51 Adam NewsomeSouth Hot-CellStandard OperationTM2 TM2 move - ITE to SHCTM2 was moved from ITE to SHC with Ta#68
target. During the move, a railing was lightly
bumped by the corner of the module - no obvious
  
  2725   Tuesday, December 03, 2024, 10:10 Frank SongITEStandard OperationTM2Ta#68module disconnectionTM2 with spent target TA#68 was disconnected 
in ITE and is being moved to SHC. All water
lines are purged for module storage.
 TM2_with_spent_target_TA#68_disconnection_in_ITE_2024-12-03.pdf 
  2724   Thursday, November 28, 2024, 16:13 Aaron TamSouth Hot-CellStandard Operation TiC#9 ParasiticsTiC#9 Parasitics Leak Check and removal from hotcell November 27th, 2024 - Aaron Tam with
Farran & David 
Parasitic Assembly Leak Check (Aaron,
 Leak_Test.jpgDissassembly.jpgSeal_Seat.jpgPEEK_seal.jpgPEEK_Seal_Closer_look.jpg 
  2723   Tuesday, November 26, 2024, 14:52 Aaron TamSouth Hot-CellStandard OperationTM3UCx#47 UC#47 LP-SIS w/ gas Install November 26, 2024 - Aaron Tam 
Handover from TPO and inspection

    Completed
    Some
 6x 
ELOG V2.9.2-2455